# Sorbonne Université, CNRS, 4 Place Jussieu, 75005 Paris, France

### v1

@@ I Introduction
It is well known that $R+\alpha R^{2}$ gravity Starobinsky:1980te provides an inflationary model that lies at the heart of the observations of CMB temperature anisotropy spectrum Planck . It is therefore legitimate to ask if and how this simple model can be generated as an effective field theory out of a fundamental theory of quantum gravity, such as string theory. There are several questions that can be posed.
@@
    $V(\phi)=\frac{1}{8{\alpha}}\left(1-e^{-\sqrt{\frac{2}{3}}\phi}\right)^{2}\,,$ (1)
    where $\phi$ is the scalaron (or inflaton in the context of inflation). The potential exponentially approaches a constant at large field values and has a minimum at $\phi=0$ with zero vacuum energy. The inflaton mass at the minimum is $m_{\phi}=1/\sqrt{6\alpha}M_{P}$. This potential is the basis of the Starobinsky model of inflation.  [^1] Field values will always be normalized to the reduced Planck mass with $M_{P}=2.4\times 10^{18}$ GeV.
@@
  • An additional puzzle is more specific to supersymmetric formulations of the Starobinsky model described by no-scale supergravity Cecotti:1987sa ; eno6 ; Ellis:2013nxa ; kl ; Ellis:2020lnc . In these constructions of the Starobinsky potential, a second (complex) scalar field must be included in addition to the inflaton (which is of course also complex), Cecotti:1987sa ; Ellis:2013nxa . Then, for successful inflation, the other three scalar degrees of freedom must be stabilized. If for example, we associate the inflaton with the real part of the ‘volume modulus’ $T$ (as in Cecotti:1987sa ), in addition to the pseudoscalar, the second field, denoted here as $C$ must have a positive mass squared with $\langle C\rangle=0$. However, in minimal constructions, $m_{C}^{2}<0$ during inflation.
@@
Our aim therefore is to identify in the string spectrum a scalar that shares very similar properties to the scalaron, having controllable quantum corrections, with the additional degrees of freedom stabilized. We will attempt to provide answers to all of the above questions.
@@ II.1 The Starobinsky model
For pedagogical reasons, we first briefly review the classical equivalence between the Starobinsky model $R+\alpha R^{2}$ and a scalar-tensor theory with a particular scalar potential. Therefore we consider the action
$S\;=\;\int[d^{4}x]\sqrt{-g}\left(\frac{1}{2}R+\frac{\alpha}{2}R^{2}\right)\,.$ (3)
To transform the action to the Einstein frame, we first introduce a Lagrange multiplier $\Phi$ identifying the scalar curvature $R$ with a scalar field $\chi$:
$S=\int[d^{4}x]\left\{{1\over 2}R+\Phi(R-\chi)+\frac{\alpha}{2}\chi^{2}\right\}\,,$ (4)
where the brackets in the integration measure include the density factor $\sqrt{|\det g|}$ and in our conventions the signature of the metric is mostly positive. Again, we remind the reader that we work in reduced Planck mass units. We proceed by integrating over $\chi$ which has a Gaussian dependence:
$S=\int[d^{4}x]\left\{{1\over 2}(1+2\Phi)R-{1\over 2\alpha}\Phi^{2}\right\}$ (5)
@@
${\tilde{g}}_{\mu\nu}\;=\;e^{2\Omega}g_{\mu\nu}\;=\;\left(1+2\Phi\right)g_{\mu\nu}\,,$ (6)
@@
$S\;=\;\int[d^{4}x]\left[\frac{1}{2}{\tilde{R}}-\frac{3}{(1+2\Phi)^{2}}\partial^{\mu}\Phi\partial_{\mu}\Phi-\frac{\Phi^{2}}{2{\alpha}(1+2\Phi)^{2}}\right]\,.$ (7)
@@
$S=\int[d^{4}x]\left\{{1\over 2}R-{1\over 2}(\partial\phi)^{2}-{1\over 8\alpha}\left(1-e^{-{\sqrt{2/3}\,\phi}}\right)^{2}\right\}\,,$ (8)
@@
Note that in all steps above from (4) to (8), no constraint on the parameter $\alpha$ was used and the above result is valid also if $\alpha$ is field dependent.
@@
$S_{\text{string}}=\int[d^{4}x]\left\{{1\over 2}R[g]-(\partial\varphi)^{2}-{1\over 2}(\partial\phi)^{2}-{e^{2\varphi}\over 8\alpha}\left(1-e^{-{\sqrt{2/3}\,\phi}}\right)^{2}\right\}\,,$ (10)
where the parameter $\alpha$ can now be treated as a numerical constant. Reversing the manipulations we have used to pass from (4) to (8), one rewrite (10) in a geometric form:
$S_{\text{string}}=\int[d^{4}x]\left\{{1\over 2}R[g]-(\partial\varphi)^{2}+\frac{\alpha}{2}e^{-2\varphi}R^{2}\right\}\,.$ (11)
Note that the dilaton exponent is consistent with a tree-level $R^{2}$ which is (globally) scale invariant and does not change under metric rescalings (modulo four-derivative interactions).
@@ II.2 Supersymmetrization
The supersymmetrization of $R+\alpha R^{2}$ has been performed in Cecotti:1987sa ; Ferrara:2013wka ; Antoniadis:2014oya . One might naively expect to be able to describe it in the context of ordinary $N=1$ supergravity coupled to a chiral multiplet corresponding to a super-Lagrange multiplier, following a procedure similar to the one described above in the bosonic theory. It turns out however, that one needs to introduce two Lagrange multipliers chiral superfields. [^2] This is related to comment made earlier regarding the construction of the Starobinsky potential in no scale supergravity requiring (at least) two chiral superfields Ellis:2013nxa . The reason is the the scalar curvature appears in the upper component of a chiral superfield $\cal R$ wesbag and therefore ${\cal R}^{2}$ does not contain $R^{2}$. The latter appears in a D-term of ${\cal R}\bar{\cal R}$ whose linearization requires two super-Lagrange multipliers. The result is an $N=1$ supergravity theory coupled to two chiral multiplets, $T$, and $C$ with a Kähler potential of the no-scale type Cremmer:1983bf ; Ellis:1983ei ; Lahanas:1986uc and superpotential given by:
@@
where $M^{2}=3/(2\alpha)$  [^3] Parallels between the $R+\alpha R^{2}$ theory and no-scale supergravity were discussed in DLT ; DGKLP ; eno9 ; Ellis:2020lnc ; Ema:2024sit . The scalar component of the superfield $T$ corresponds to the Starobinsky inflaton with $T_{\rm R}=\text{Re}T=\frac{1}{2}e^{\sqrt{2/3}\phi}$, with the same scalar potential (1) when $T_{\rm I}=\text{Im}T=C=0$. The superfield $C$ contains the goldstino during inflation, as can be seen from the linear term in the superpotential (12).
@@
where the factors $(T+{\bar{T}})/3$ and $(T+{\bar{T}})^{2}/3$ come from the normalization of the kinetic terms. It follows that $T_{\rm I}$ has a large positive mass-squared and we can safely set $T_{\rm I}=0$. In this case, $C$ becomes tachyonic for $T_{\rm R}>(\sqrt{2}+1)/2\simeq 1.2$. This corresponds to $\phi\simeq 1.1<\phi_{*}$, making the inflationary trajectory unstable to any fluctuations in the $C$ (or ${\bar{C}}$) direction. One way to cure this instability during inflation is to modify the Kähler potential by adding a quartic $(C{\bar{C}})^{2}$ in the argument of the logarithm Ellis:1984bs ; Ellis:2013nxa ; Ellis:2015kqa ; Ellis:2018zya ; Kaneta:2019yjn ; Ellis:2020lnc ; Ellis:2020xmk . But this comes at the expense of losing the geometric interpretation of the (supersymmetric) Starobinsky model. We discuss an alternative solution to the stabilization $C$ in Section III.
@@ II.3 An explicit string model
Along the direction $y_{I}=z=0$ we can redefine $y_{R}=\tanh(\phi/\sqrt{6})$, where $y\equiv y_{R}+iy_{I}$ and $\phi$ is the canonically normalized inflaton. We then obtain the same potential as in Eq. (1) with $M^{2}=3/(2\alpha)$ (as it must, since we simply performed a field redefinition). Now, $\phi_{*}=5.35$, corresponds to $y_{R}\simeq 0.975$.
@@
The potential (21) is actually part of a larger class of models known as $\alpha$-Starobinsky or no-scale attractor models Ellis:2013nxa ; KLR ; Kallosh:2014rga ; Roest:2015qya ; Scalisi:2015qga ; ENOV3 ; Ellis:2020xmk ; Saini:2024mun . In the context of no-scale supergravity, we can replace the coefficient of the logarithm in Eqs. (12) and (16), $3\to 3\alpha$. Indeed, it was shown in Ellis:2013nxa , that for a potential $\propto(1-e^{-\sqrt{2/3\alpha}\phi})^{2}$ leads to a value of $n_{s}$ independent of $\alpha$, whereas $r\propto\alpha$. It is clear then that the $R+R^{2}$ construction requires $\alpha=1$, whereas the model described by Eq. (18) corresponds to $\alpha=2/3$ and accounts for the difference in the predicted value of $r$.
@@ IV Generalization in a wider class of no-scale supergravity models
With this ansätz, both $T_{\rm I}=0$ and $\phi=0$ are extrema. For $T_{\rm I}=0$, to be a minimum, $f_{T}^{2}-ff_{TT}$ must be positive ($f_{T}=df/dT$). This is of course true for the Starobinsky model derived from Eq. (12) with $f(T)=T-1/2$. For the sgoldstino, the direction $C=0$ is also an extremum, and the condition for a minimum (along $T_{\rm I}=0$) is $f^{2}-4Tff_{T}+2T^{2}f_{T}^{2}>0$. However, during inflation or wherever the potential is nearly flat, $Tf_{T}\simeq f$ in which case
@@
Note that this model is equivalent to a ‘dual’ model obtained by replacing in the superpotential (36), $f(T)$ with ${\tilde{f}}(T)=(\lambda T)^{2}f({1\over\lambda T})$, with $\lambda$ an arbitrary constant. This can be easily seen using the fact that $T\to 1/(\lambda^{2}T)$ and $C\to C/(\lambda T)$ amounts to a Kähler transformation $K\to K+3\ln(\lambda T)+3\ln(\lambda{\bar{T}})$. In the case of $f(T)=T-1/2$, choosing $\lambda=2$, one finds that ${\tilde{f}}(T)=2T(T-1/2)$ and the canonical field is $T_{\rm R}=\frac{1}{2}e^{-\sqrt{2/3}\phi}$.
@@
will reduce to Eq. (38) along $T_{\rm I}=C=0$. The relation between $T$ and the canonical $\phi$ is altered to $T=\frac{1}{2}e^{\sqrt{2/3\alpha}\phi}$. For the choice $f=T-1/2$, the potential becomes
@@
The condition for $T_{\rm I}$ to be a minimum is now altered. It is $f_{T}^{2}-ff_{TT}+\frac{3}{2}(\alpha-1)f^{2}/T^{2}>0$. For the ($\alpha$)-Starobinsky choice $f=T-1/2$, $T_{\rm I}=0$ is always a minimum if $\alpha\geq 1$. For $11/3<\alpha<$ 1, there is a tachyonic instability appearing at small $T_{\rm R}<1/4$ corresponding to $\phi_{\rm tach}<-\ln 2/\sqrt{2}$. However, this instability is actually never reached as during its oscillations $|\phi_{\rm tach}|>\phi_{\rm end}$, where the latter is the inflaton field value when exponential expansion ends and oscillations begin. For $\alpha=1$, $\phi_{\rm end}\simeq 0.62$, whereas for $\alpha=1/3$, $\phi_{\rm end}\simeq 0.49$ and the amplitude of oscillations decreases with each oscillation as $|\phi|^{2}\propto a^{-3}$, where $a$ is the cosmological scale factor. For $\alpha<1/3$, there is another tachyonic instability which occurs at large $T$ and hence may occur during inflation. The condition for $C=0$ to be a minimum is unchanged from the previous set of models, and during inflation, the mass-squared of $C$ is still given by (39).  [^4] For the dual model, $T_{\rm I}=0$ is always a minimum so long as $\alpha\geq 1/3$. For smaller $\alpha<1/3$, there is again a tachyonic instability, which now occurs at small $T_{\rm R}$, which in this case is during inflation.
@@
which reduces to (38) modulo the factor of $(S+{\bar{S}})^{-1}$ for $C=0$. Although the extremal condition for $T_{I}=0$ is unchanged, the condition for a minimum at $C=0$ is now $05f^{2}-8Tff_{T}+4T^{2}f_{T}^{2}>$ 0. But now during inflation, using $Tf_{T}\simeq f$ , instead of (39), we have

### v2

@@ I Introduction
It is well known that $R+{\tilde{\alpha}}R^{2}$ gravity Starobinsky:1980te provides an inflationary model that lies at the heart of the observations of CMB temperature anisotropy spectrum Planck . It is therefore legitimate to ask if and how this simple model can be generated as an effective field theory out of a fundamental theory of quantum gravity, such as string theory. There are several questions that can be posed.
@@
    $V(\phi)=\frac{1}{8{{\tilde{\alpha}}}}\left(1-e^{-\sqrt{\frac{2}{3}}\phi}\right)^{2}\,$, (1)
    where $\phi$ is the scalaron (or inflaton in the context of inflation). The potential exponentially approaches a constant at large field values and has a minimum at $\phi=0$ with zero vacuum energy. The inflaton mass at the minimum is $m_{\phi}=1/\sqrt{6{\tilde{\alpha}}}M_{P}$. This potential is the basis of the Starobinsky model of inflation.  [^1] Field values will always be normalized to the reduced Planck mass with $M_{P}=2.4\times 10^{18}$ GeV.
@@
  • An additional puzzle is more specific to supersymmetric formulations of the Starobinsky model described by no-scale supergravity Cecotti:1987sa ; eno6 ; Ellis:2013nxa ; kl ; Ellis:2020lnc . In these constructions of the Starobinsky potential, a second (complex) scalar field must be included in addition to the inflaton (which is of course also complex), Cecotti:1987sa ; Ellis:2013nxa . Then, for successful inflation, the other three scalar degrees of freedom must be stabilized. If for example, we associate the inflaton with the real part of the ‘volume modulus’ $T$ (as in Cecotti:1987sa ), in addition to the pseudoscalar, the second field, denoted here as $C$ must have a positive mass squared with $\langle C\rangle=0$. However, in minimal constructions, $m_{C}^{2}<0$ during inflation. Actually $C$ is the goldstino superfield of the spontaneously broken supersymmetry during inflation.
@@
In the context of string theory, an $R^{2}$ term cannot appear as such, since it contains an extra degree of freedom, which should be part of the free string spectrum. Our aim therefore is to identify in the string spectrum a scalar that shares very similar properties to the Starobinsky scalaron. Moreover, here we restrict to $N=1$ supersymmetric compactifications which also provide a framework for interesting particle physics phenomenology incorporating a supersymmetric extension of the Standard Model. One should therefore identify within the massless string spectrum two chiral multiplets that generate a scalar potential similar to $R+R^{2}$ supergravity, having controllable quantum corrections, with the additional degrees of freedom (beyond the scalaron) stabilized. We will then attempt to provide answers to all of the above questions.
@@ II.1 The Starobinsky model
For pedagogical reasons, we first briefly review the classical equivalence between the Starobinsky model $R+{\tilde{\alpha}}R^{2}$ and a scalar-tensor theory with a particular scalar potential. Therefore we consider the action
$S\;=\;\int d^{4}x\,\sqrt{-{\tilde{g}}}\left(\frac{1}{2}R[\tilde{g}]+\frac{{\tilde{\alpha}}}{2}R[\tilde{g}]^{2}\right)\,$, (3)
where $\tilde{g}$ is the ‘Jordan’ frame metric (to be clarified below). To transform the action to the Einstein frame, we first introduce a Lagrange multiplier $\Phi$ identifying the scalar curvature $\tilde{R}$ with a scalar field $\chi$:
$S=\int[d^{4}x]\left\{{1\over 2}\tilde{R}+\Phi(\tilde{R}-\chi)+\frac{{\tilde{\alpha}}}{2}\chi^{2}\right\}\,$, (4)
where the brackets in the integration measure include the density factor $\sqrt{|\det\tilde{g}|}$ and in our conventions the signature of the metric is mostly positive. Again, we remind the reader that we work in reduced Planck mass units. We proceed by integrating over $\chi$ which has a Gaussian dependence:
$S=\int[d^{4}x]\left\{{1\over 2}(1+2\Phi)\tilde{R}-{1\over 2{\tilde{\alpha}}}\Phi^{2}\right\}$ (5)
@@
$g_{\mu\nu}\;=\;e^{2\Omega}\tilde{g}_{\mu\nu}\;=\;\left(1+2\Phi\right)\tilde{g}_{\mu\nu}\,$, (6)
@@
$S\;=\;\int[d^{4}x]\left[\frac{1}{2}R[g]-\frac{3}{(1+2\Phi)^{2}}\partial^{\mu}\Phi\partial_{\mu}\Phi-\frac{\Phi^{2}}{2{{\tilde{\alpha}}}(1+2\Phi)^{2}}\right]\,.$ (7)
@@
$S=\int[d^{4}x]\left\{{1\over 2}R-{1\over 2}(\partial\phi)^{2}-{1\over 8{\tilde{\alpha}}}\left(1-e^{-{\sqrt{2/3}\,\phi}}\right)^{2}\right\}\,$, (8)
@@
Note that in all steps above from (4) to (8), no constraint on the parameter $\tilde{\alpha}$} was used and the above result is valid also if $\tilde{\alpha}$} is field dependent.
@@
$S_{\text{string}}=\int[d^{4}x]\left\{{1\over 2}R[g]-(\partial\varphi)^{2}-{1\over 2}(\partial\phi)^{2}-{e^{2\varphi}\over 8{\tilde{\alpha}}}\left(1-e^{-{\sqrt{2/3}\,\phi}}\right)^{2}\right\}\,$, (10)
where the parameter $\tilde{\alpha}$} can now be treated as a numerical constant. Reversing the manipulations we have used to pass from (4) to (8), one rewrite (10) in a geometric form:
$S_{\text{string}}=\int[d^{4}x]\left\{{1\over 2}R[\tilde{g}]-(\tilde{\partial}\varphi)^{2}+\frac{{\tilde{\alpha}}}{2}e^{-2\varphi}\left[\tilde{R}-2(\tilde{\partial}\varphi)^{2}\right]^{2}\right\}\,.$ (11)
The interested reader can verify (10) starting from (11) by following exactly the steps (4) to (8) and replacing $\tilde{R}$ by $\left[\tilde{R}-2(\tilde{\partial}\varphi)^{2}\right]$ in (4), (5) and $R$ by $\left[R-2({\partial}\varphi)^{2}\right]$ in (7), (8) recalling that both $R$ and $(\partial\varphi)^{2}$ transform with the same conformal factor $e^{2\Omega}$ which is still given by Eq. (6).
Note that the dilaton exponent in (11) is consistent with a tree-level $R^{2}$ which is (globally) scale invariant and does not change under metric rescalings (modulo four-derivative interactions of the dilaton and metric).
@@ II.2 Supersymmetrization
[…]
@@
[…]
@@
where the factors $(T+{\bar{T}})/3$ and $(T+{\bar{T}})^{2}/3$ come from the normalization of the kinetic terms. It follows that $T_{\rm I}$ has a large positive mass-squared and we can safely set $T_{\rm I}=0$. In this case, $C$ becomes tachyonic for $T_{\rm R}>(\sqrt{2}+1)/2\simeq 1.2$. This corresponds to $\phi\simeq 1.1<\phi_{*}$, making the inflationary trajectory unstable to any fluctuations in the $C$ (or $\bar{C}$}) direction. One way to cure this instability during inflation is to modify the Kähler potential by adding a quartic $(C{\bar{C}})^{2}$ in the argument of the logarithm Ellis:1984bs ; Ellis:2013nxa ; Ellis:2015kqa ; Ellis:2018zya ; Kaneta:2019yjn ; Ellis:2020lnc ; Ellis:2020xmk . But this comes at the expense of losing the geometric interpretation of the (supersymmetric) Starobinsky model. Another approach is to parametrize the model dependent stabilization of the sgoldstino by integrating it out to an effective field theory obtained by imposing a non-linear supersymmetric nilpotent constraint $C^{2}=0$ which eliminates the scalar component in terms of the goldstino bilinear Antoniadis:2014oya . Here, we choose to not follow this approach since we would like to keep $N=1$ linear supersymmetry at low energies, as we described in the introduction. Instead, we discuss an alternative solution to the stabilization of $C$ in Section III.
@@ II.3 An explicit string model
[…]
@@
The potential (21) is actually part of a larger class of models known as $\alpha$-Starobinsky or no-scale attractor models Ellis:2013nxa ; KLR ; Kallosh:2014rga ; Roest:2015qya ; Scalisi:2015qga ; ENOV3 ; Ellis:2020xmk ; Saini:2024mun . In the context of no-scale supergravity, we can replace the coefficient of the logarithm in Eqs. (12) and (16), $3\to 3\alpha$. Indeed, it was shown in Ellis:2013nxa , that for a potential $\propto(1-e^{-\sqrt{2/3\alpha}\phi})^{2}$ leads to a value of $n_{s}$ independent of $\alpha$, whereas $r\propto\alpha$. It is clear then that the $R+{\tilde{\alpha}}R^{2}$ construction requires $\alpha=1$, whereas the model described by Eq. (18) corresponds to $\alpha=2/3$ and accounts for the difference in the predicted value of $r$.
@@ IV Generalization in a wider class of no-scale supergravity models
With this ansätz, both $T_{\rm I}=0$ and $\phi=0$ are extrema. For $T_{\rm I}=0$, to be a minimum, $f_{T}^{2}-ff_{TT}$ must be positive ($f_{T}=df/dT$). This is of course true for the Starobinsky model derived from Eq. (12) with f(T)=T1/2f(T)=T-1/2. For the sgoldstino, the direction $C=0$ is also an extremum, and the condition for a minimum (along $T_{\rm I}=0$) is $f^{2}-4Tff_{T}+2T^{2}f_{T}^{2}>0$. However, during inflation or wherever the potential is nearly flat, $Tf_{T}\simeq f$ in which case
@@
Note that this model is equivalent to a ‘dual’ model obtained by replacing in the superpotential (36), $f(T)$ with ${\tilde{f}}(T)=(\lambda T)^{2}f({1\over\lambda T})$, with $\lambda$ an arbitrary constant. This can be easily seen using the fact that $T\to 1/(\lambda^{2}T)$ and $C\to C/(\lambda T)$ amounts to a Kähler transformation $K\to K+3\ln(\lambda T)+3\ln(\lambda{\bar{T}})$. In the case of f(T)=T1/2f(T)=T-1/2, choosing $\lambda=2$, one finds that ${\tilde{f}}(T)=2T(T-1/2)$ and the canonical field is $T_{\rm R}=\frac{1}{2}e^{-\sqrt{2/3}\phi}$.
@@
will reduce to Eq. (38) along $T_{\rm I}=C=0$. The relation between $T$ and the canonical $\phi$ is altered to $T=\frac{1}{2}e^{\sqrt{2/3\alpha}\phi}$. For the choice f=T1/2f=T-1/2, the potential becomes
@@
The condition for $T_{\rm I}$ to be a minimum is now altered. It is $f_{T}^{2}-ff_{TT}+\frac{3}{2}(\alpha-1)f^{2}/T^{2}>0$. For the ($\alpha$)-Starobinsky choice f=T1/2f=T-1/2, $T_{\rm I}=0$ is always a minimum if $\alpha\geq 1$. For $1/3<\alpha<1$, there is a tachyonic instability appearing at small $T_{\rm R}<1/4$ corresponding to $\phi_{\rm tach}<-\ln 2/\sqrt{2}$. However, this instability is actually never reached as during its oscillations $|\phi_{\rm tach}|>\phi_{\rm end}$, where the latter is the inflaton field value when exponential expansion ends and oscillations begin. For $\alpha=1$, $\phi_{\rm end}\simeq 0.62$, whereas for $\alpha=1/3$, $\phi_{\rm end}\simeq 0.49$ and the amplitude of oscillations decreases with each oscillation as $|\phi|^{2}\propto a^{-3}$, where $a$ is the cosmological scale factor. For $\alpha<1/3$, there is another tachyonic instability which occurs at large $T$ and hence may occur during inflation. The condition for $C=0$ to be a minimum is unchanged from the previous set of models, and during inflation, the mass-squared of $C$ is still given by (39).  [^4] For the dual model, $T_{\rm I}=0$ is always a minimum so long as $\alpha\geq 1/3$. For smaller $\alpha<1/3$, there is again a tachyonic instability, which now occurs at small $T_{\rm R}$, which in this case is during inflation.
@@
which reduces to (38) modulo the factor of $(S+{\bar{S}})^{-1}$ for $C=0$. Although the extremal condition for $T_{I}=0$ is unchanged, the condition for a minimum at $C=0$ is now $5f^{2}-8Tff_{T}+4T^{2}f_{T}^{2}>0$. But now during inflation, using $Tf_{T}\simeq f$ , instead of (39), we have